OpenStudy (anonymous):

how would write this into simplified form? the third power to the square root of 24

OpenStudy (anonymous):

\[\sqrt[3]{24}\]

OpenStudy (kropot72):

\[\sqrt[3]{24}=\sqrt[3]{3\times 8}=2\sqrt[3]{3}\]

OpenStudy (kropot72):

The radical sign with the small 3 in front means the cube root of the term under the sign. In this case 24 can be factorised to give a factor (8) that has a simple cube root.

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)